A fan at STP is at .87 “ WC delivers 20,800 cfm at 1700 RPM and draws 3.9 BHP. The fan is moved to an altitude of 9500 feet ASL. Find the CFM delivered at the same RPM. What is the SP and HP at the new altitude? What was the number of lbs/ hr delivered at SL versus at altitude? How much must the fan be sped up to deliver the same mass flow? What is the new SP and BHP at the increased speed?


1
Expert's answer
2021-11-01T08:52:58-0400

Cfm=rpm*area

="1700*(87*12*9500)"

=1.6861*1010cfm

"Hp=\frac{cfm*p}{229*\mu}"

"=\frac{1.6861*10^{10}*3.9}{229*1}"

"=27.8Ghp"

"Sp=\frac{BHP*6356}{cfm*sg}"

"=\frac{3.9*6356}{1.6861*10^{10}*1}"

"=1.4702*10^{-6}"